OpenStudy (anonymous):

kay tears off labels of all 10 soup cans her mother knows there are 2 tomato and 8 vegetable she selects 4 at random. what is the probability a. exactly one of the 4 cans is tomato b. none of the 4 cans is tomato c. at least one of the 4 is tomato?

OpenStudy (anonymous):

a) 2C1/10C4

OpenStudy (anonymous):

b) 2C0/10C4

OpenStudy (anonymous):

c) (2C1 + 2C2)/10C4

OpenStudy (anonymous):

i am very confused...and i know the answers to these problems...i just dont know how to do them?

OpenStudy (anonymous):

on top of the fraction,the numerator is the no of ways o selecting 1 tomato can out of a possible 2 , and the denominator is the total no. of selections ie 10C4, selecting any 4 out of 10 cans

OpenStudy (anonymous):

get it or nt?

OpenStudy (anonymous):

no not really........the answer to the first is .8 so that is why i am confused

OpenStudy (anonymous):

cant be

OpenStudy (anonymous):

how?

OpenStudy (amistre64):

10! -------- sounds plausible? 2!(10-2)!

OpenStudy (anonymous):

check my solution...shes sayin its 0.8 the answer

OpenStudy (amistre64):

then it aint 45 lol

OpenStudy (amistre64):

2 2 2 2 ------- perchance? 10 9 8 7

OpenStudy (amistre64):

combo and perms are still new to me :)

OpenStudy (anonymous):

ha same

OpenStudy (amistre64):

id have to draw up a tree or something and count them out lol

OpenStudy (anonymous):

yeah...i am not sure if it is a combo though? def probablility

OpenStudy (anonymous):

its using combos

OpenStudy (anonymous):

m 95% sure my methods right nelly

OpenStudy (amistre64):

there are 45 ways to select 4 cans right? or did i mess that up too

OpenStudy (anonymous):

10C4

OpenStudy (amistre64):

10C4....what is that in longhand tho?

OpenStudy (amistre64):

10! --- ?? 4!6!

OpenStudy (anonymous):

10!/(6!4!)

OpenStudy (amistre64):

210?

OpenStudy (anonymous):

i got 120960?

OpenStudy (amistre64):

10 9 8 7 6 5 4 3 2 ---------------- 4 3 2 6 5 4 3 2 10 9 8 7 -------- 4 3 2 5 3 2 7 = 15(14) = 210 right?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

so we have 210 ways to do this lol

OpenStudy (anonymous):

great

OpenStudy (amistre64):

how many ways can we have 2 tcans?

OpenStudy (anonymous):

672

OpenStudy (amistre64):

or is it easier to say how many ways to pick no tcans? 8 can ways then...

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

pick one out of 2 cans is 2C1

OpenStudy (amistre64):

but that doesnt gives us a total ways for tcans does it?

OpenStudy (anonymous):

336

OpenStudy (amistre64):

if there are 8 ways to choose 4 cans 8! --- ?? 4!4!

OpenStudy (amistre64):

70?

OpenStudy (amistre64):

8765432 -------- 432.432 8765 ----- 432 725 = 14*5 = 70

OpenStudy (amistre64):

70/210 = 7/21 = 1/3; 33.33% ??

OpenStudy (amistre64):

so the chance that 1 is at least a tcan could be 66' 2/3% does that seem reasonable?

OpenStudy (amistre64):

33' 1/3% chance we got no tcans 66' 2/3% chance we got at least 1 tcan seems good to me

OpenStudy (anonymous):

nope the anwser is .8 blah

OpenStudy (amistre64):

.8 is 80% which answer is .8? a,b or c?

OpenStudy (amistre64):

10C4 is the total number of ways to choose 4 cans.... 8C4 is the total number of ways to choose no tcans.... casue we removed them from the equation.... % = part/total % = 8c4/10c4 for no tcans 100- no tcan% = at least 1 tcan.... is my reasoning

OpenStudy (anonymous):

a

OpenStudy (amistre64):

whats the answers to b and c :)

OpenStudy (anonymous):

b .333333 c. .66666

OpenStudy (amistre64):

yay!! i was right :)

OpenStudy (amistre64):

exactly one of the 4 cans is tomato lets see if we take away one tcan; just for the sake of argument; we can do this then

OpenStudy (amistre64):

9! ---- = 3.3.2.7 = 9(14) = 126 ways to possible ways to 4!5! to choose at least 1 tcan

OpenStudy (anonymous):

i am still very confused how to do the first as i get 122/210 = .53333

OpenStudy (amistre64):

126 ways for 1 possible tcan and we still have 70 ways for no tcans..; i wonder

OpenStudy (amistre64):

126-70 would give us the number of ways for 1 tcan right? 126 70 ---- 56 ways that at least 1 tcan is in it... i hope lol

OpenStudy (amistre64):

56/210 . --------- 105/28.0000 ......not gonna be .8 lol